# C.V. Raman Avenue, Bengaluru 560012, India

### 18 Dec 2023 v1

@@ Abstract
Building upon recent research in spin systems with non-local interactions, this study investigates operator growth using the Krylov complexity in different non-local versions of the Ising model. We find that the non-locality results in a faster scrambling of the operator to all sites. While the saturation value of Krylov complexity of local integrable and local chaotic theories differ by a significant margin, this difference is much suppressed when non-local terms are introduced in both regimes. This results from the faster scrambling of information in the presence of non-locality. In addition, we investigate the behavior of level statistics and spectral form factor as probes of quantum chaos to study the integrability breaking due to non-local interactions. Our numerics indicate that even in the non-local case, Krylov complexity can distinguish between different underlying theories as well as different degrees of non-locality.
1 Introduction
@@
The study of operator growth using Krylov complexity (also referred to as ‘K-complexity’) has emerged as a powerful framework for investigating the behavior of quantum systems and their underlying dynamics PhysRevX.9.041017 ; barbonEvolutionOperatorComplexity2019a ; rabinoviciOperatorComplexityJourney2021b . K-complexity is a measure of the delocalization of a local initial operator evolving under Heisenberg evolution with respect to the Hamiltonian PhysRevX.9.041017 ; barbonEvolutionOperatorComplexity2019a ; rabinoviciOperatorComplexityJourney2021b ; PhysRevB.102.085137 ; jianComplexityGrowthOperators2021 . It is conjectured to grow at most exponentially generically non-integrable systems PhysRevX.9.041017 . This exponential growth of K-complexity can be used to extract the Lyapunov exponent PhysRevX.9.041017 ; barbonEvolutionOperatorComplexity2019a , establishing a connection with out-of-time-ordered-correlators (OTOC) maldacenaBoundChaos2016 ; PhysRevResearch.2.043234 . Further studies have shown a relation between K-complexity and chaos in context of various models such as Ising models rabinovici_krylov_2022 ; rabinoviciKrylovLocalizationSuppression2022 , Sachdev-Ye-Kitaev (SYK) models jianComplexityGrowthOperators2021 ; bhattacharjeeKrylovComplexityLarge2023 ; heQuantumChaosScrambling2022 , quantum field theories caputaOperatorGrowth2d2021 ; khetrapalChaosOperatorGrowth2023 ; kunduStateDependenceKrylov2023 ; camargoKrylovComplexityFree2023 ; avdoshkinKrylovComplexityQuantum2022 ; Erdmenger:2023wjg , many-body localization system 10.21468/SciPostPhys.13.2.037 , and open quantum systems bhattacharyaOperatorGrowthKrylov2022b ; bhattacharjeeOperatorGrowthOpen2023 ; bhattacharyaKrylovComplexityOpen2023d ; PhysRevResearch.5.033085 .
Among different systems, of particular interest are the so-called fast scramblers for which scrambling time is logarithmic in the number of degrees of freedom. The examples include black holes Yasuhiro_Sekino_2008 ; lashkariFastScramblingConjecture2013 ; maldacenaBoundChaos2016 , conjectured to be the fastest scramblers in nature, SYK PhysRevLett.70.3339 and other related holographic models. The previous studies PhysRevA.94.040302 ; PhysRevA.99.051803 ; PhysRevLett.123.130601 ; belyansky2020minimal suggest that local chaotic dynamics along with non-local interactions are sufficient to give rise to the phenomena termed as “fast scrambling”. These developments have motivated the study of quantum information scrambling in non-local systems PhysRevResearch.2.043399 ; wanisch2022quantum as well as experimental proposal for probing fast scramblers with simpler models PhysRevLett.123.130601 ; belyansky2020minimal ; PhysRevResearch.2.043399 ; PhysRevB.107.L081103 .
@@
In this study, we characterize the growth of K-complexity in non-local systems. It has been established that K-complexity can distinguish between the integrable and chaotic regimes in local systems. To be precise, the complexity growth rate and saturation value are known to be significantly higher for chaotic systems as compared to the integrable ones rabinoviciOperatorComplexityJourney2021b ; rabinovici_krylov_2022 ; rabinoviciKrylovLocalizationSuppression2022 . Here, we ask the question of how these characteristics get modified once non-local interactions are turned on within these regimes. While chaotic systems with non-local interactions can exhibit fast scrambling properties, the non-local terms in the otherwise integrable cases are also expected to accelerate the operator growth and therefore begs some investigation as to whether the notions of local integrability and chaos remain distinguishable in the presence of non-locality.
@@
Apart from K-Complexity, we also study the two widely used measures to probe chaos: Level Statistics (Appendix A.1) and Spectral Form Factor (Appendix A.2). The symmetry-reduced level statistics is known to exhibit Poisson distribution for integrable regimes, while in chaotic regimes, it shows the Wigner-Dyson distribution 10.1007/978-3-0348-8266-8_36 ; PhysRevLett.52.1 ; PhysRevLett.110.084101 . On the other hand, the spectral form factor (SFF) is found to exhibit an explicit dip-ramp-plateau behaviour for chaotic evolution in contrast to the integrable regime where higher amount of fluctuations wash out the dip-ramp-plateau behaviour cotlerBlackHolesRandom2017 ; PhysRevD.95.126008 ; winerSpectralFormFactor2022 . [^2] This might vary from model to model. In some cases, an effective ramp appears, but it is not as evidently clear as it is for chaotic ones. We study these quantities to compare their behaviour in the presence of non-local terms in the Hamiltonian and analyse how the introduction of non-locality changes their nature.
Our results indicate that with increasing degrees of non-locality, the integrable evolution becomes similar to chaotic ones. We study the three diagnostics for characterizing chaos for various models which indicates that non locality apparently makes all regimes chaotic. However, a closer look at the plots reveals that non-local terms with local integrable operator evolution can still be distinguished from non-local terms with local chaotic operator evolution using K-complexity. Moreover, we also observe that the increase in non-locality is captured by the initial growth rate of the K-complexity. To have a clearer understanding of how non-locality changes behavior of Krylov complexity, we also study “mixed field” spin chains wanisch2022quantum , where we vary the degree of non-locality and study how it increases the slope of the Lanczos ascent.
@@
The rest of the paper is structured as follows. In section 2, we briefly review the notion of K-complexity, and it’s characteristics for local Hamiltonians, discussing both integrable and non-integrable regimes. In section 3, we introduce several non-local models and discuss distinguished features present in the K-complexity profiles. Section 3.1 consists of the numerical findings for non-locality introduced to an otherwise local integrable and chaotic version of the transverse field Ising model. In sections 3.2 and 3.3, we present the models and numerics when we deal with varying degrees of non-locality by tuning a parameter in transverse field Ising model and XXZ spin chain respectively. Finally, we conclude in section 4 discussing the implications of our results. In Appendix we study the remaining probes, namely the level statistics and the spectral form Factor for both local and non-local models to be able to compare their behavior with Krylov complexity.
2 Review for local Hamiltonians
@@
where $g$ and $h$ are the coupling parameters. When $h=0$, the Hamiltonian remains integrable for all values of $g$, as it can be mapped to the free-fermionic model sachdev_2011 . However, when both $g$ and $h$ are nonzero, the system departs from integrability. To investigate the integrable regime, we set $g=1$ and $h=0$, while for the chaotic regime, we choose $g=-1.05$ and $h=0.5$ PhysRevLett.106.050405 . The model has a parity symmetry for all values of coupling parameters, and additional $Z$-reflection symmetry when $h=0$.
2.1 Operator dynamics in Krylov basis
We begin with a brief review of the operator dynamics in Krylov spacePhysRevX.9.041017 ; rabinoviciOperatorComplexityJourney2021b ; barbonEvolutionOperatorComplexity2019a . Consider the evolution of an operator with seed $\mathcal{O}_{0}$ at $t=0$ under the time-independent Hamiltonian $H$. At any time $t$, the time-evolved operator $\mathcal{O}(t)$ under Heisenberg evolution can be written as
@@ 2.1 Operator dynamics in Krylov basis
$\mathcal{H}_{\mathcal{O}}=\text{span}\{\mathcal{L}^{n}\mathcal{O}\}^{\infty}_{n=0}=\text{span}\{\mathcal{O},[H,\mathcal{O}],[H,[H,\mathcal{O}]],\ldots\}$ (3)
Once the Krylov space is obtained, an orthonormal basis is then formed using a choice of inner product on the operator space. This is achieved using the Lanczos algorithm, an instance of the Gram-Schmidt process PhysRevX.9.041017 ; baiTemplatesSolutionAlgebraic1987 . For our purpose, we will use the infinite-temperature inner product, also known as Frobenius inner product :
@@
where K is dimension of Krylov space, constrained to $11\leq K\leq D^{2}-D+$ 1 rabinoviciOperatorComplexityJourney2021b . Eq. (5) can be recast into the following form,
@@
We can also define the complexity operator hornedal_ultimate_2022 ,
@@
where $i$ is chosen somewhere close to the center of the chain. Fig. 1 shows the Lanczos sequence $b_{n}$ and K-complexity for TFIM Hamiltonian (1) computed for $L=7$ for fixed parity $P=+1$ sector for the operator $\mathcal{O}=S^{z}_{4}$. The study of Lanczos sequence and Krylov complexity in integrable and chaotic limits have extensively been done in previous studies for different systems rabinovici_krylov_2022 ; iizukaKrylovComplexityIP2023 ; bhattacharyyaOperatorGrowthKrylov2023 ; camargoSpectralKrylovComplexity2023 . The Lanczos coefficient features sublinear growth in the integrable limit while linear growth in the chaotic limit, followed by saturation, and the descent, while K-complexity transitions from exponential growth at very early time to linear increase followed by saturation rabinoviciOperatorComplexityJourney2021b ; rabinovici_krylov_2022 . The saturation value of K-complexity is significantly large for local chaotic evolution as compared to the local integrable one. The suppressed complexity saturation of integrable cases results from higher fluctuations in Lanczos coefficients during the descent phase. This fluctuating descent period is dubbed as the Krylov localization.
3 Non-local models and numerical findings:
In this section, we will study operator growth in non-local systems. We compare the results with the local counterpart to demonstrate the effect of non-locality on these quantities. For these Hamiltonians, it is, in general, expected that due to non-local terms, the Liouvillian will also be non-local. Now, since the notion of integrability is believed to be closely related to locality PhysRevResearch.2.043399 ; wanisch2022quantum , the evolution of the operator under the non-local Liouvillian is expected to show chaotic behavior even in the case when the local part of the Hamiltonian is integrable. Therefore it is expected that the non-local terms in the Hamiltonian will result in the loss of its integrability.
@@ 3 Non-local models and numerical findings:
3.1 Non-local Transverse Field Ising Model
We consider the fast scrambling spin-1/2 Hamiltonian introduced in belyansky2020minimal ,
@@ 3.1 Non-local Transverse Field Ising Model
where $\sigma^{z}_{i}$ is the Pauli $z$ operator acting on site $i$ and $\mathcal{H}_{\text{local}}$ is Hamiltonian with only local interactions. In our study, we will consider the local Hamiltonian as a one-dimensional transverse-field Ising model with open boundary conditions introduced in Sec. 2. At this point, it is important to emphasize the terminology used throughout this literature. We will later observe from the results of operator growth as well as level-statistics and spectral form factor (discussed in the appendices A.1 and A.2), that non-locality can break the integrability of the system. While referring to a system whose local part of the Hamiltonian in Eq. (14) is integrable, we will call it non-local integrable Hamiltonian even though non-locality breaks integrability, and the model is no longer integrable.
Figure 2: Comparison between Lanczos-coefficients for local and non-local TFIM Hamiltonian with non-local parameter $\gamma=0.5$, computed for $N=7$ spins in the $P=+1$ sector for Integrable and Chaotic limits of local Hamiltonian. Top row: complete spectrum of Lanczos coefficients, Bottom row: Initial lanczos coefficients.
@@
  • Lanczos sequence The initial growth of Lanczos coefficients is faster for the non-local cases as compared to local cases (in both integrable and chaotic limits). The peak value of the Lanczos sequence is more than that of the corresponding local case. The overall Lanczos profile for the non-local integrable case resembles the form of the local chaotic case. The fluctuations in the decaying part of the Lanczos sequence are also small for the non-local integrable case which results in a high saturation value of K-complexity.
@@
3.2 Transverse Mixed Field Ising Model
@@
where $J^{\alpha}_{ij}$ is interaction strength between spins at position $i$ and $j$ which assume to follow powerlaw $1/\kappa\cdot J/|i-j|^{\alpha}$. Throughout our study, we take $J=1$, and $\kappa=1$. In previous studies wanisch2022quantum , the information scrambling, using out-of-time-order correlators (OTOCs) as a probe, has been studied in such a model with a variation of non-local parameter $\alpha$. It’s shown that for $\alpha>2$, the dynamics effectively resemble local dynamics, while for smaller exponents, the dynamics become non-local. Here, we will confine ourselves to studying the initial Lanczos sequence for the model with varying exponent $\alpha$.
@@
In Fig. 4, we showed the initial Lanczos coefficients $b_{n}$ for $L=13$ with varying exponent $\alpha\in[0.1,0.5,1.0,1.5,2.0,2.5]$. The initial operator is chosen to be $S^{z}_{7}$. In both limits integrable and chaotic, we find the increasing slope and saturation value in initial growth with the decrease in exponent $\alpha$ and therefore increase in non-locality. This ensures that the initial (pre-scrambling) growth rate of the Krylov complexity always increases in the presence of non-locality (See Fig. 5). In Figure 6, we showed the linear growth rate calculated using linear fit ($b_{n}=\delta n+c$) in initial Lanczos coefficients.
Figure 6: The linear growth rate $\delta$ plotted for different non-local exponent $\alpha$ values for Lanczos coefficients in Fig. 4.
3.3 Mixed field XXZ Hamiltonian

### 11 Mar 2024 v2

# Krylov complexity for non-local spin chains

@@ Abstract
Abstract Building upon recent research in spin systems with non-local interactions, this study investigates operator growth using the Krylov complexity in different non-local versions of the Ising model. We find that the non-locality results in a faster scrambling of the operator to all sites. While the saturation value of Krylov complexity of local integrable and local chaotic theories differ by a significant margin, this difference is much suppressed when non-local terms are introduced in both regimes. This results from the faster scrambling of information in the presence of non-locality. In addition, we investigate the behavior of level statistics and spectral form factor as probes of quantum chaos to study the integrability breaking due to non-local interactions. Our numerics indicate that in the non-local case, late time saturation of Krylov complexity distinguishes between different underlying theories, while the early time complexity growth distinguishes different degrees of non-locality.
DOI: 10.1103/PhysRevD.109.066010
I   Introduction
@@
The study of operator growth using Krylov complexity (also referred to as ‘K-complexity’) has emerged as a powerful framework for investigating the behavior of quantum systems and their underlying dynamics [1, 2, 3]. K-complexity is a measure of the delocalization of a local initial operator evolving under Heisenberg evolution with respect to the Hamiltonian [1, 2, 3, 4, 5]. It is conjectured to grow at most exponentially generically non-integrable systems [1]. This exponential growth of K-complexity can be used to extract the Lyapunov exponent [1, 2], establishing a connection with out-of-time-ordered-correlators (OTOC) [6, 7]. Further studies have shown a relation between K-complexity and chaos in context of various models such as Ising models [8, 9], Sachdev-Ye-Kitaev (SYK) models [5, 10, 11], quantum field theories [12, 13, 14, 15, 16, 17], many-body localization system [18, 19], and open quantum systems [20, 21, 22, 23, 24]. Among different systems, of particular interest are the so-called fast scramblers for which scrambling time is logarithmic in the number of degrees of freedom. The examples include black holes [25, 26, 6], conjectured to be the fastest scramblers in nature, SYK [27] and other related holographic models. The previous studies [28, 29, 30, 31] suggest that local chaotic dynamics along with non-local interactions are sufficient to give rise to the phenomena termed as “fast scrambling". These developments have motivated the study of quantum information scrambling in non-local systems [32, 33] as well as experimental proposal for probing fast scramblers with simpler models [30, 31, 32, 34].
@@
In this study, we characterize the growth of K-complexity in non-local systems. It has been established that K-complexity can distinguish between the integrable and chaotic regimes in local systems. To be precise, the complexity growth rate and saturation value are known to be significantly higher for chaotic systems as compared to the integrable ones [3, 8, 9]. Here, we ask the question of how these characteristics get modified once non-local interactions are turned on within these regimes. While chaotic systems with non-local interactions can exhibit fast scrambling properties, the non-local terms in the otherwise integrable cases are also expected to accelerate the operator growth and therefore begs some investigation as to whether the notions of local integrability and chaos remain distinguishable in the presence of non-locality.
@@
Apart from K-Complexity, we also study the two widely used measures to probe chaos: Level Statistics (Appendix A.1) and Spectral Form Factor (Appendix A.2). The symmetry-reduced level statistics is known to exhibit Poisson distribution for integrable regimes, while in chaotic regimes, it shows the Wigner-Dyson distribution [35, 36, 37]. On the other hand, the spectral form factor (SFF) is found to exhibit an explicit dip-ramp-plateau behaviour for chaotic evolution in contrast to the integrable regime where higher amount of fluctuations wash out the dip-ramp-plateau behaviour [38, 39, 40]. [^1] This might vary from model to model. In some cases, an effective ramp appears, but it is not as evidently clear as it is for chaotic ones. We study these quantities to compare their behaviour in the presence of non-local terms in the Hamiltonian and analyse how the introduction of non-locality changes their nature.
@@
Our results indicate that with increasing degrees of non-locality, the integrable evolution becomes similar to chaotic ones. We study the three diagnostics for characterizing chaos for various models which indicates that non locality apparently makes all regimes chaotic. However, a closer look at the plots reveals that non-local terms with local integrable operator evolution can still be distinguished from non-local terms with local chaotic operator evolution using K-complexity. Moreover, we also observe that the increase in non-locality is captured by the initial growth rate of the K-complexity. To have a clearer understanding of how non-locality changes behavior of Krylov complexity, we also study “mixed field" spin chains [33], where we vary the degree of non-locality and study how it increases the slope of the Lanczos ascent. The rest of the paper is structured as follows. In section II, we briefly review the notion of K-complexity, and it’s characteristics for local Hamiltonians, discussing both integrable and non-integrable regimes. In section III, we introduce several non-local models and discuss distinguished features present in the K-complexity profiles. Section III.1 consists of the numerical findings for non-locality introduced to an otherwise local integrable and chaotic version of the transverse field Ising model. In sections III.2 and III.3, we present the models and numerics when we deal with varying degrees of non-locality by tuning a parameter in transverse field Ising model and XXZ spin chain respectively. Finally, we conclude in section IV discussing the implications of our results. In Appendix we study the remaining probes, namely the level statistics and the spectral form Factor for both local and non-local models to be able to compare their behavior with Krylov complexity.
II   Review for local Hamiltonians
@@
where $g$ and $h$ are the coupling parameters. When $h=0$, the Hamiltonian remains integrable for all values of $g$, as it can be mapped to the free-fermionic model [41]. However, when both $g$ and $h$ are nonzero, the system departs from integrability. To investigate the integrable regime, we set $g=1$ and $h=0$, while for the chaotic regime, we choose $g=-1.05$ and $h=0.5$ [42]. The model has a parity symmetry for all values of coupling parameters, and additional $Z$-reflection symmetry when $h=0$.
II.1   Operator dynamics in Krylov basis
We begin with a brief review of the operator dynamics in Krylov space[1, 3, 2]. Consider the evolution of an operator with seed $\mathcal{O}_{0}$ at $t=0$ under the time-independent Hamiltonian $H$. At any time $t$, the time-evolved operator $\mathcal{O}(t)$ under Heisenberg evolution can be written as
@@ II.1   Operator dynamics in Krylov basis
$\begin{split}\mathcal{H}_{\mathcal{O}}&=\text{span}\{\mathcal{L}^{n}\mathcal{O}\}^{\infty}_{n=0}\\ &=\text{span}\{\mathcal{O},[H,\mathcal{O}],[H,[H,\mathcal{O}]],\ldots\}\end{split}$ (3)
Once the Krylov space is obtained, an orthonormal basis is then formed using a choice of inner product on the operator space. This is achieved using the Lanczos algorithm, an instance of the Gram-Schmidt process [1, 43]. For our purpose, we will use the infinite-temperature inner product, also known as Frobenius inner product :
@@
where $K$ is dimension of Krylov space, constrained to $11\leq K\leq D^{2}-D+$ 1 [3]. Eq. (5) can be recast into the following form,
@@
We can also define the complexity operator [44],
@@
where $i$ is chosen somewhere close to the center of the chain. [^2] The operator is chosen so that it respects the parity symmetry (defined as in [45]) and keeps the computation within the chosen sector. The results deduced in this work remains unchanged even if $i$ is varied in Eq. (13). However, as discussed in [8], the behavior of K-complexity is controlled by the statistics of the Hamiltonian spectrum and the structure of operator under consideration. For example, if the chosen operator has spread in all the sites from the beginning, the behaviour of Krylov complexity might change as there is no scope for scrambling. Therefore, while the results can differ for an arbitrary seed operator, the conclusions remain invariant for any operator that respects the symmetries (parity here) of the Hamiltonian and has support in small number of sites to begin with so that it has enough scope to spread and scramble in the rest of the sites. Fig. 1 shows the Lanczos sequence $b_{n}$ and K-complexity for TFIM Hamiltonian (1) computed for $L=7$ for fixed parity $P=+1$ sector for the operator $\mathcal{O}=S^{z}_{4}$. The study of Lanczos sequence and Krylov complexity in integrable and chaotic limits have extensively been done in previous studies for different systems [8, 46, 47, 48]. The Lanczos coefficient features sublinear growth in the integrable limit while linear growth in the chaotic limit, followed by saturation, and the descent, while K-complexity transitions from exponential growth at very early time to linear increase followed by saturation [3, 8]. The saturation value of K-complexity is significantly large for local chaotic evolution as compared to the local integrable one. The suppressed complexity saturation of integrable cases results from higher fluctuations in Lanczos coefficients during the descent phase. This fluctuating descent period is dubbed as the Krylov localization.
III   Non-local models and numerical findings:
In this section, we will study operator growth in non-local systems. We compare the results with the local counterpart to demonstrate the effect of non-locality on these quantities. For these Hamiltonians, it is, in general, expected that due to non-local terms, the Liouvillian will also be non-local. Now, since the notion of integrability is believed to be closely related to locality [32, 33], the evolution of the operator under the non-local Liouvillian is expected to show chaotic behavior even in the case when the local part of the Hamiltonian is integrable. Therefore it is expected that the non-local terms in the Hamiltonian will result in the loss of its integrability.
@@ III   Non-local models and numerical findings:
III.1   Non-local Transverse Field Ising Model
We consider the fast scrambling spin-1/2 Hamiltonian introduced in [31],
@@ III.1   Non-local Transverse Field Ising Model
where $\sigma^{z}_{i}$ is the Pauli $z$ operator acting on site $i$ and $\mathcal{H}_{\text{local}}$ is Hamiltonian with only local interactions. In our study, we will consider the local Hamiltonian as a one-dimensional transverse-field Ising model with open boundary conditions introduced in Sec. II. At this point, it is important to emphasize the terminology used throughout this literature. We will later observe from the results of operator growth as well as level-statistics and spectral form factor (discussed in the appendices A.1 and A.2), that non-locality can break the integrability of the system. While referring to a system whose local part of the Hamiltonian in Eq. (14) is integrable, we will call it non-local integrable Hamiltonian even though non-locality breaks integrability, and the model is no longer integrable.
Figure 2: Comparison between Lanczos-coefficients for local and non-local TFIM Hamiltonian with non-local parameter $\gamma=0.5$, computed for $L=7$ spins in the $P=+1$ sector for Integrable and Chaotic limits of local Hamiltonian. Top row: complete spectrum of Lanczos coefficients, Bottom row: Initial lanczos coefficients.
@@
Figure 4: Comparison between initial Lanczos-coefficients for local and non-local TFIM Hamiltonian with non-local parameter $\gamma=0.5$, computed for $L=12$ spins in the $P=+1$ sector for Integrable and Chaotic limits of local Hamiltonian. The Lanczos coefficients exhibit faster growth with non-local terms, especially evident in larger system sizes.
@@
  • Lanczos sequence The initial growth of Lanczos coefficients is faster for the non-local cases as compared to local cases (in both integrable and chaotic limits). While this accelerated growth is not prominently observable in smaller system sizes, it becomes distinctly apparent in larger system sizes as shown in Figure 4. Note that the growth of the primary few Lanczos coefficients correspond to the pre-scrambling behaviour. Hence a faster growth of primary Lanczos coefficients result in an higher exponent for the exponentially growing complexity before scrambling time. The peak value of the Lanczos sequence is more than that of the corresponding local case. The overall Lanczos profile for the non-local integrable case resembles the form of the local chaotic case. The fluctuations in the decaying part of the Lanczos sequence are also small for the non-local integrable case which results in a high saturation value of K-complexity.
@@
III.2   Transverse Mixed Field Ising Model
@@
where $J^{\alpha}_{ij}$ is interaction strength between spins at position $i$ and $j$ which assume to follow powerlaw $1/\kappa\cdot J/|i-j|^{\alpha}$. Throughout our study, we take $J=1$, and $\kappa=1$. In previous studies [33], the information scrambling, using out-of-time-order correlators (OTOCs) as a probe, has been studied in such a model with a variation of non-local parameter $\alpha$. It’s shown that for $\alpha>2$, the dynamics effectively resemble local dynamics, while for smaller exponents, the dynamics become non-local. Here, we will confine ourselves to studying the initial Lanczos sequence for the model with varying exponent $\alpha$.
@@
In Fig. 5, we showed the initial Lanczos coefficients $b_{n}$ for $L=13$ with varying exponent $\alpha\in\{0.1,0.5,1.0,1.5,2.0,2.5\}$. The initial operator is chosen to be $S^{z}_{7}$. In both limits integrable and chaotic, we find the increasing slope and saturation value in initial growth with the decrease in exponent $\alpha$ and therefore increase in non-locality. This ensures that the initial (pre-scrambling) growth rate of the Krylov complexity always increases in the presence of non-locality (See Fig. 6). In Figure 7, we showed the initial growth rate calculated using fit ,$b_{n}=\delta\frac{n}{ln[n]}+c$ (for $1$D systems, as conjectured in [1]) in initial Lanczos coefficients.
Figure 7: The growth rate $\delta$ plotted for different non-local exponent $\alpha$ values for Lanczos coefficients in Fig. 5.
Figure 8: The complete lanczos-sequence is calculated for mixed-field TFIM for $L=7$ spins in the $P=+1$ sector. The comaprison between the coefficients is shown for integrable and chaotic values with Left: $\alpha=0.5$ Right: $\alpha=1$. The inset shows that the fluctuations for integrable parameter value (blue) are larger compared to chaotic parameter value (red). These larger fluctuation in large $b_{n}$ leads to suppression of K-complexity for integrable parameter values even for non-local model.
Figure 9: Late time Krylov complexity for the Transverse Mixed Field Ising Model Hamiltonian computed for $L=7$ spins with open boundary conditions for fixed parity $P=+1$ in (a) Integrable (b) Chaotic limits with varying non-local exponent $\alpha$ values.
The saturation value of the late-time Krylov Complexity is controlled by the amount of fluctuations in the Lanczos coefficients as reported in [9, 22]. Fluctuations are more for integrable parameter values and this phenomenon is dubbed as the Krylov localization. We notice that although non-local terms in the Hamiltonian increase the initial growth in Lanczos coefficients for both integrable and chaotic parameter values, the fluctuations in later Lanczos coefficients is still comparatively more for the integrable parameter values. This is shown in Figure 8 for two choices of the non-local exponent $\alpha$ of the mixed TFIM model. As a result, the late time saturation value of Krylov complexity in Figure 9 distinctly differentiates the nature of the underlying integrable and chaotic theories even in the non-local models. In Figure 9, the left figure shows the complexity saturation for various non-local $+$ integrable cases while the same is shown for non-local $+$ chaotic cases for various $\alpha$. While the late-time saturation value indeed distinguishes whether the underlying theory is integrable or chaotic, it remains insensitive to the magnitude of the non-local exponent $\alpha$. Nevertheless, as we noted earlier, the pre-scrambling growth of both Lanczos coefficients as well as the Krylov complexity serves as a clear indicator distinguishing between the non-local behaviors associated with different values of the exponent (Figure 5).
III.3   Mixed field XXZ Hamiltonian